OpenStudy (anonymous):

if the sum of three consecutive number is 75 find the number can you show me how you got the answer Thanks I think I got it now

OpenStudy (anonymous):

let the consecutiv numbers be x-1 , x and x+1 now proceed with the problem

OpenStudy (anonymous):

@bhhoward direct answers are not encouraged here

OpenStudy (anonymous):

sum will be x-1 + x + x +1 = 3x which is given as 75 then x = 75/3 = 25

OpenStudy (anonymous):

then x - 1 = 25 - 1 = 24 is one numbr and x + 1 = 25 + 1 = 26 is another and we already got 25

OpenStudy (anonymous):

@bhhoward If you like to work the things out in your way, then it might not be the good place to start with and if you dont know the "official " way to do the things here, you would have studied the code of conduct http://openstudy.com/code-of-conduct .

OpenStudy (ja1):

@bhhoward I am in complete agreement with @SandeepReddy , OpenStudy is a place to learn and teach, not one to give direct answers, try and guide the user to the answer by asking him/her questions and involving them in the solving process. If you don't know exactly how to set it up then don't be afraid to ask for help, we are open to collaboration and anyone can pitch in to help the user! So let's set our differences apart here and help :) http://openstudy.com/code-of-conduct